# Leveraging Machine Learning for High-Dimensional Option Pricing within the Uncertain Volatility Model

### v1

@@ Abstract
This paper explores the application of Machine Learning techniques for pricing high-dimensional options within the framework of the Uncertain Volatility Model (UVM). The UVM is a robust framework that accounts for the inherent unpredictability of market volatility by setting upper and lower bounds on volatility and the correlation among underlying assets. By leveraging historical data and extreme values of estimated volatilities and correlations, the model establishes a confidence interval for future volatility and correlations, thus providing a more realistic approach to option pricing. By integrating advanced Machine Learning algorithms, we aim to enhance the accuracy and efficiency of option pricing under the UVM, especially when the option price depends on a large number of variables, such as in basket or path-dependent options. Our approach evolves backward in time, dynamically selecting at each time step the most expensive volatility and correlation for each market state. Specifically, it identifies the particular values of volatility and correlation that maximize the expected option value at the next time step. This is achieved through the use of Gaussian Process regression, the computation of expectations via a single step of a multidimensional tree and the Sequential Quadratic Programming optimization algorithm. The numerical results demonstrate that the proposed approach can significantly improve the precision of option pricing and risk management strategies compared with methods already in the literature, particularly in high-dimensional contexts.
Keywords: UVM, option pricing, GPR, Tree method, SQP.
@@ 1 Introduction
The Uncertain Volatility Model represents a significant advancement in financial modeling, specifically for the pricing and hedging of derivative securities in markets characterized by uncertain volatility. Developed initially by Avellaneda et al. [1], the UVM addresses the limitations of traditional models like the Black-Scholes, which assume constant volatility over the life of an option. Instead, the UVM assumes that volatility fluctuates within a known range, providing a more realistic framework for financial markets where volatility is inherently unpredictable.
@@
The core idea of the UVM is to set upper and lower bounds on volatility, denoted as $\sigma^{\min}$ and $\sigma^{\max}$. These bounds can be inferred from historical data or extreme values of implied volatilities from liquid options. The model uses these bounds to establish a confidence interval within which the future volatility is expected to lie. This approach acknowledges the inherent uncertainty in predicting future volatility and provides a robust framework for pricing options under this uncertainty [1, 8].
@@
One of the seminal contributions of Avellaneda et al. [1] was the derivation of the Black-Scholes-Barenblatt (BSB) equation, a non-linear partial differential equation (PDE) that extends the classical Black-Scholes equation by incorporating the variable volatility bounds. This PDE dynamically selects the "pricing" volatility from the two extreme values based on the convexity of the value function. For instance, if the second derivative (Gamma) of the option price is positive, the model uses $\sigma^{\max}$, otherwise, it uses $\sigma^{\min}$.
@@
The UVM framework is particularly effective for superhedging strategies, where traders aim to hedge against the worst-case scenarios of volatility. By ensuring that the option pricing incorporates the highest potential volatility, traders can guarantee non-negative profit and loss outcomes even when actual realized volatility deviates significantly from the expected values. This is especially useful in markets with high volatility or during periods of market stress, as pointed out by Martini and Jacquier [8].
@@
In addition to its theoretical foundations, the UVM has practical implications for risk management. Pooley et al. [9] provided a detailed numerical analysis of the UVM, highlighting its advantages in constructing efficient hedges and managing derivatives positions in environments with uncertain volatility. Their work emphasized the importance of diversification in managing volatility risk and provided algorithms for implementing the UVM in real-world trading systems.
@@
Several extensions and practical applications of the UVM have been explored in subsequent research. For instance, Windcliff et al. [13] use methods based on partial differential equations (PDEs) to evaluate cliquet options in the context of the uncertain volatility model. The paper explores how to solve the PDEs associated with these financial instruments by comparing different numerical and grid construction techniques. It is shown that the PDE-based approach is effective in capturing the complexities of the UVM model, providing a more accurate evaluation of cliquet options than other traditional methods. Martini and Jacquier [8] elaborated on the superhedging strategies within the UVM framework, discussing the implications for derivative desks where risk aversion is a critical factor. They demonstrated that the UVM could be applied to a range of option types, including those with non-convex payoffs, by employing more sophisticated numerical methods such as trinomial trees and finite-difference schemes.
@@
More recently, Guyon and Labordère [6] introduce two innovative Monte Carlo methods to price financial derivatives under the UVM. These methods address the challenges posed by high-dimensionality, where traditional finite-difference schemes fall short. Moreover, the paper considers a path-dependent option, which is particularly useful in the context of the UVM. This is beneficial because path-dependent options can capture the effects of the asset’s historical prices, providing a more accurate valuation under uncertainty, and aligning well with the UVM’s focus on modeling worst-case scenarios. Despite one may appreciate the innovative approach presented in the paper, it is worth noting that the numerical results are occasionally suboptimal, and the authors do not extend their numerical analysis to derivatives with more than two underlyings.
@@
In this paper, we propose an innovative approach to the problem of derivative evaluation within the UVM. Specifically, the proposed method is based on the GPR-Tree pricing algorithm, which we introduced in a previous paper [5] for the valuation of options dependent on multiple underlyings. Consequently, we refer to this method as GTU, which stands for GPR-Tree method for UVM. Our algorithm involves calculating the price in the UVM by working backward in time and determining, for selected market states, the maximum price by leveraging an enhanced tree method and a suitable optimization algorithm. This calculation accounts for varying volatilities and correlations between the underlyings. The obtained prices for specific typical market states are then generalized to all possible market conditions by using Gaussian Process Regression (GPR), a powerful machine learning technique for multidimensional regressions from scattered data.
@@
The proposed method not only serves as a viable alternative to existing methods for handling problems in lower dimensions but also demonstrates exceptional capability in valuing derivatives in high-dimensional settings, including scenarios with dozens of underlyings. To the best of our knowledge, this is the first instance where such a high-dimensional problem has been addressed effectively. Moreover, the proposed method is particularly effective for problems where the high dimension arises not from the number of underlyings, but from the impact of the entire process history on the derivative’s final value. This is especially relevant for path-dependent options, which our algorithm can evaluate proficiently.
The paper includes the results of several numerical experiments, which show that the GTU algorithm is reliable and fast. In particular, we observe that the results returned by GTU are very stable and very close to the benchmark prices when these are available.
@@
The remainder of the paper is organized as follows. Section 2 introduces the UVM and the pricing principles relevant to this context. Section 3 reviews the fundamentals of the GPR-Tree algorithm. Section 4 demonstrates the evolution of the GPR-Tree algorithm into the proposed GTU method. Section 5 provides detailed numerical results from key experiments. Finally, Section 6 concludes the paper.
@@ 2 The Multidimensional Uncertain Volatility Model
where $\mathbf{S}_{0}=(s_{0}^{1},\dots,s_{0}^{d})\in\mathbb{R}_{+}^{d}$ is the initial spot price vector, $r$ is the constant risk-free interest rate, $\mathbf{\eta}=(\eta_{1},\dots,\eta_{d})^{\top}$ is the vector of dividend yields, $\mathbf{\sigma}=(\sigma_{1},\dots,\sigma_{d})^{\top}$ is the vector of volatilities, $\mathbf{W}$ is a $d$-dimensional correlated Wiener processes, and $\rho_{i,j}$ is the instantaneous correlation coefficient between $W^{i}$ and $W^{j}$. We also term $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ the natural filtration generated by the $\mathbf{W}$.
@@
$dS_{t}^{i}=S_{t}^{i}(\left(r-\eta_{i}\right)dt+\sigma_{i}\Sigma_{i}d\mathbf{B}_{t}),$ (2.2)
where $\mathbf{B}$ is a $d$-dimensional uncorrelated Wiener processes and $\Sigma_{i}$ is the $i$-th row of the matrix
$\Sigma=\Sigma\left(\rho_{1,2},\dots,\rho_{d-1,d}\right)$ (2.3)
defined as a square root of the correlation matrix $\Gamma$, given by
@@
$\mathbf{S}_{t}=\mathbf{S}_{0}\circ\exp\left(\left(r-\mathbf{\eta}-\frac{\sigma^{2}}{2}\right)\Delta t+\sigma\circ\left(\Sigma\mathbf{B}_{t}\right)\right)$ (2.5)
with $\mathbf{v}\circ\mathbf{w}$ the Hadamard product between vectors $\mathbf{v}$ and $\mathbf{w}$, that is the element-wise product, and the exponential function in (2.5) is applyed to each component of the vector $\left(r-\frac{\sigma^{2}}{2}\right)\Delta t+\sigma\circ\left(\Sigma\mathbf{B}_{t}\right)$.
Instead of assuming a single constant volatility, the UVM considers a range of possible volatilities. Specifically, the vector of volatilities $\mathbf{\sigma}$ is not known, but it can change over time, as a function of the time and of the underlyings, so we write $\mathbf{\sigma}=\sigma\left(t,\mathbf{S}_{t}\right)=\left(\sigma_{1}\left(t,\mathbf{S}_{t}\right),\dots,\sigma_{d}\left(t,\mathbf{S}_{t}\right)\right)^{\top}$. Moreover, the explicit form of $\sigma\left(t,\mathbf{S}_{t}\right)$ is not specified in the model, but we only know that, for each $i=1,\dots,d$ there exist two bounds for $\sigma_{i}$, that is
@@
Such a degree of uncertainty can be extended to correlations between Wiener processes. Specifically $\rho_{i,j}=\rho_{i,j}\left(t,\mathbf{S}_{t}\right)$ may change over time based on the underlying values, but the particular law is not a model datum. Instead, we only know that, for each couple of values $i=1,\dots,d-1$ and $j=i+1,\dots,d$, there exist two bounds for $\rho_{i,j}$, that is
@@
$P=P\left(t,\mathbf{S}_{t}\right)=\left(\sigma_{1}\left(t,\mathbf{S}_{t}\right),\dots,\sigma_{d,t}\left(t,\mathbf{S}_{t}\right),\rho_{1,2}\left(t,\mathbf{S}_{t}\right),\dots,\rho_{d-1,d}\left(t,\mathbf{S}_{t}\right)\right)^{\top}$
the vector of all uncertain model parameters as a function of $t$ and $\mathbf{S}_{t}$, and $\mathcal{A}$ the set of all admissible values $P$, that is, which satisfy relations (2.6), (2.7), and (2.9).
In the UVM, the pricing of derivatives is approached as an optimization problem due to the uncertainty in the volatility and correlation of the underlying asset. This model aims to find the worst-case scenario within this range to ensure robust and conservative pricing. Specifically, we seek to maximize the expected payoff of the derivative over all admissible scenarios for volatilities and correlation, that is over all $P$ in $\mathcal{A}$. Let $V\left(t,\mathbf{S}_{t}\right)$ represent the value of the derivative at time $t$ and asset price $\mathbf{S}_{t}$ in the UVM. The pricing problem is then given by:
$V\left(t,\mathbf{S}_{t}\right)=\sup_{P\in\mathcal{A}}\mathbb{E}\left[e^{-r(T-t)}\Psi\left(\mathbf{S}_{T}\right)\mid\mathbf{S}_{t}\right],$ (2.10)
where $\Psi$ is the payoff function at maturity $T$, the dynamics of $\mathbf{S}_{t}$ is given by (2.8) for $P\in\mathcal{A}$, and $\mathbb{E}$ denotes the expectation under the risk-neutral measure $\mathbb{Q}$.
This optimization ensures that the derivative’s price accounts for the most adverse volatility-correlation scenario, providing a conservative estimate that is robust to volatility fluctuations. By solving this maximization problem, we obtain a price that safeguards against the uncertainty in the market, making it a reliable valuation method in uncertain volatility conditions.
3 The GPR-Tree Method
The proposed GTU method is based on the GPR-Tree algorithm. Here we give some details about this latter algorithm and we refer the interested reader for further details to [5].
The GPR-Tree algorithm exploits Gaussian Gaussian Process Regression and a multimensional binomial tree to compute the price of multidimensional American option.

### v2

@@ Abstract
This paper explores the application of Machine Learning techniques for pricing high-dimensional options within the framework of the Uncertain Volatility Model (UVM). The UVM is a robust framework that accounts for the inherent unpredictability of market volatility by setting upper and lower bounds on volatility and the correlation among underlying assets. By leveraging historical data and extreme values of estimated volatilities and correlations, the model establishes a confidence interval for future volatility and correlations, thus providing a more realistic approach to option pricing. By integrating advanced Machine Learning algorithms, we aim to enhance the accuracy and efficiency of option pricing under the UVM, especially when the option price depends on a large number of variables, such as in basket or path-dependent options. In this paper, we consider two approaches based on Machine Learning. The first one, termed GTU, evolves backward in time, dynamically selecting at each time step the most expensive volatility and correlation for each market state. Specifically, it identifies the particular values of volatility and correlation that maximize the expected option value at the next time step, and therefore, an optimization problem must be solved. This is achieved through the use of Gaussian Process regression, the computation of expectations via a single step of a multidimensional tree and the Sequential Quadratic Programming optimization algorithm. The second approach, referred to as NNU, leverages neural networks and frames pricing in the UVM as a control problem. Specifically, we train a neural network to determine the most adverse volatility and correlation for each simulated market state, generated via random simulations. The option price is then obtained through Monte Carlo simulations, which are performed using the values for the uncertain parameters provided by the neural network. The numerical results demonstrate that the proposed approaches can significantly improve the precision of option pricing and risk management strategies compared with methods already in the literature, particularly in high-dimensional contexts.
Keywords: Machine Learning, Gaussian Process Regression, Neural Networks, Option Pricing, Uncertain Volatility Model, Stochastic Control.
@@ 1 Introduction
The Uncertain Volatility Model represents a significant advancement in financial modeling, specifically for the pricing and hedging of derivative securities in markets characterized by uncertain volatility. Developed initially by [Avellaneda et al., 1995], the UVM addresses the limitations of traditional models like the Black-Scholes, which assume constant volatility over the life of an option. Instead, the UVM assumes that volatility fluctuates within a known range, providing a more realistic framework for financial markets where volatility is inherently unpredictable.
@@
The core idea of the UVM is to set lower and upper bounds on volatility, denoted as $\sigma^{\min}$ and $\sigma^{\max}$. These bounds can be inferred from historical data or extreme values of implied volatilities from liquid options. The model uses these bounds to establish a confidence interval within which the future volatility is expected to lie. This approach acknowledges the inherent uncertainty in predicting future volatility and provides a robust framework for pricing options under this uncertainty, as discussed in [Avellaneda et al., 1995] and [Martini and Jacquier, 2008].
@@
One of the seminal contributions of [Avellaneda et al., 1995] was the derivation of the Black-Scholes-Barenblatt (BSB) equation, a non-linear partial differential equation (PDE) that extends the classical Black-Scholes equation by incorporating the variable volatility bounds. This PDE dynamically selects the "pricing" volatility from the two extreme values based on the convexity of the value function. For instance, if the second derivative (Gamma) of the option price is positive, the model uses $\sigma^{\max}$, otherwise, it uses $\sigma^{\min}$.
@@
The UVM framework is particularly effective for superhedging strategies, where traders aim to hedge against the worst-case scenarios of volatility. By ensuring that the option pricing incorporates the highest potential volatility, traders can guarantee non-negative profit and loss outcomes even when actual realized volatility deviates significantly from the expected values. This is especially useful in markets with high volatility or during periods of market stress, as pointed out in [Martini and Jacquier, 2008].
@@
In addition to its theoretical foundations, the UVM has practical implications for risk management. [Pooley et al., 2003] provided a detailed numerical analysis of the UVM, highlighting its advantages in constructing efficient hedges and managing derivatives positions in environments with uncertain volatility. Their work emphasized the importance of diversification in managing volatility risk and provided algorithms for implementing the UVM in real-world trading systems.
@@
Several extensions and practical applications of the UVM have been explored in subsequent research. For instance, [Windcliff et al., 2006] use methods based on PDEs to evaluate cliquet options in the context of the uncertain volatility model. The paper explores how to solve the PDEs associated with these financial instruments by comparing different numerical and grid construction techniques. It is shown that the PDE-based approach is effective in capturing the complexities of the UVM model, providing a more accurate evaluation of cliquet options than other traditional methods. [Martini and Jacquier, 2008] elaborate on the superhedging strategies within the UVM framework, discussing the implications for derivative desks where risk aversion is a critical factor. Moreover, they demonstrate that the UVM could be applied to a range of option types, including those with non-convex payoffs, by employing more sophisticated numerical methods such as trinomial trees and finite-difference schemes.
@@
Subsequently, [Guyon and Henry-Labordère, 2011] introduce two innovative Monte Carlo methods for pricing financial derivatives under the UVM. These methods address the challenges posed by high-dimensionality, where traditional finite-difference schemes fall short. Moreover, the paper considers a path-dependent option, which is particularly useful in the context of the UVM. This is beneficial because path-dependent options can capture the effects of the asset’s historical prices, providing a more accurate valuation under uncertainty, and aligning well with the UVM’s focus on modeling worst-case scenarios. Despite one may appreciate the innovative approach presented in the paper, it is worth noting that the numerical results are occasionally suboptimal, and the authors do not extend their numerical analysis to derivatives with more than two underlyings.
More recently, the literature on parameter uncertainty has gained renewed momentum. In this regard, [Cohen and Tegnér, 2019] explore European option pricing under parameter uncertainty in stochastic volatility models by framing the uncertainty as a control problem and employing backward stochastic differential equations. Similarly, [Lütkebohmert et al., 2022] tackle the challenges posed by uncertain model parameters in generalized affine processes by leveraging a dynamic programming approach enhanced with deep neural networks. Furthermore, [Akhtari et al., 2023] develop a generalized Feynman-Kac formula to address volatility uncertainty, extending its applicability to settings involving a discounting term in the associated PDE.
Over the last decade, a growing body of research has demonstrated that modern Machine Learning (ML) techniques can substantially enhance the valuation of derivative securities. Early contributions focus on Gaussian Process Regression (GPR), a non-parametric Bayesian method well-suited to interpolating scattered data in high-dimensional spaces. For instance, [Ludkovski, 2018a] evaluate Bermudan options by regressing continuation values on GPR surrogates, while [Goudenège et al., 2020] introduce three GPR-based schemes to price American basket options within multi-asset Black-Scholes dynamics.
Beyond GPR, neural networks (NNs) have become a cornerstone of data-driven option pricing. By casting valuation as a high-dimensional function-approximation task, both feed-forward and convolutional architectures have been trained to learn the mapping from underlying price paths to discounted payoffs, thereby mitigating the curse of dimensionality that hampers classical Monte-Carlo methods, as demonstrated in [E et al., 2017] and [Han et al., 2018]. More recently, Deep Hedging – a method introduced in [Buehler et al., 2019] – combines semi-recurrent networks with reinforcement-learning objectives to learn optimal hedging and pricing strategies directly from market data, obviating explicit model calibration. Within the regression-based Longstaff-Schwartz framework, [Lapeyre and Lelong, 2021] replace the conventional least-squares step by a neural estimator, and [Becker et al., 2019, Becker et al., 2020, Becker et al., 2021] develop deep-learning algorithms that address a broad class of American-style valuation and stochastic control problems.
Collectively, these advances illustrate how ML models can furnish accurate approximations of continuation values, accelerate pricing in large state spaces, and enable data-driven hedging policies-thereby extending the frontier of tractable problems in derivative valuation.
@@
In this paper, we propose two innovative approaches to the problem of high-dimensional derivative evaluation under parameter uncertainty, specifically within the UVM. In particular, the first one is based on the GPR-Tree pricing algorithm, which we introduced in a previous paper [Goudenège et al., 2020] for the valuation of options dependent on multiple underlyings. Consequently, we refer to this method as GTU, which stands for GPR-Tree method for UVM. This algorithm involves calculating the price in the UVM by working backward in time and determining, for selected market states, the maximum price by leveraging an enhanced tree method and a suitable optimization algorithm. This calculation accounts for varying volatilities and correlations between the underlyings. The obtained prices for specific typical market states are then generalized to all possible market conditions by using GPR, a powerful ML technique for multidimensional regressions from scattered data.
We emphasize that the GTU algorithm solves the valuation problem backward in time by determining the optimal values of the uncertain parameters at each step and for each state of the underlying assets. This approach is analogous to the pricing of American options, where the key is to determine the optimal exercise strategy through backward induction. In this regard, we refer the reader to the seminal papers by [Longstaff and Schwartz, 2001, Haugh and Kogan, 2004, Andersen and Broadie, 2004, Nadarajah et al., 2017].
The second methodology we introduce builds upon a deep learning framework in which a neural network is trained to manage parameter uncertainty. Consequently, we refer to this method as NNU, which stands for neural network method for UVM. At each time step, the network dynamically selects key diffusion parameters – i.e. volatilities and correlations – with the objective of optimising a prescribed terminal criterion. This strategy combines Monte Carlo simulation, differentiable control, and penalisation techniques to ensure tractable and robust training, even under structural constraints like positive definiteness of correlation matrices.
It is important to emphasize that the neural network employed in this approach does not operate through backward induction. Instead, it is trained in a forward-global manner: the same model is applied at every time step, and its input includes the current time and the simulated asset values up to that point. As a result, the network learns a globally consistent control policy over the entire time horizon, without relying on dynamic programming or backward recursion.
@@
The proposed methods not only serve as viable alternatives to existing classic methods for handling problems in lower dimensions but also demonstrate exceptional capability in valuing derivatives in high-dimensional settings, including scenarios with dozens of underlyings. To the best of our knowledge, this is the first instance where such a uncertain-parameters high-dimensional problem has been addressed effectively. Moreover, the proposed methods are particularly effective for problems where the high dimension arises not from the number of underlyings, but from the impact of the entire process history on the derivative’s final value. This is especially relevant for path-dependent options, which our algorithms can evaluate proficiently, such as the Call Sharpe option which is discussed in this manuscript.
The paper presents several numerical experiments demonstrating that the proposed algorithms are both accurate and computationally efficient. In particular, the results exhibit high stability and remain consistently close to benchmark prices whenever such references are available.
@@
The remainder of the paper is organized as follows. Section 2 introduces the UVM and the pricing principles relevant to this context. Section 3 reviews the fundamentals of the GPR-Tree algorithm and demonstrates the evolution of the GPR-Tree algorithm into the proposed GTU method. Section 4 outlines the foundational concepts of NNs and presents how they are employed into the proposed NNU methodology. Section 5 provides detailed numerical results from key experiments. Finally, Section 6 concludes the paper.
@@ 2 The Multidimensional Uncertain Volatility Model
where $\mathbf{S}_{0}=(s_{0}^{1},\dots,s_{0}^{d})\in\mathbb{R}_{+}^{d}$ is the initial spot price vector, $r$ is the constant risk-free interest rate, $\boldsymbol{\eta}=(\eta_{1},\dots,\eta_{d})^{\top}$ is the vector of dividend yields, $\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{d})^{\top}$ is the vector of volatilities, $\mathbf{W}=(\mathbf{W}_{t})_{t\in[0,T]}$ is a $d$-dimensional correlated Wiener processes, and $\rho_{i,j}$ is the instantaneous correlation coefficient between $W^{i}$ and $W^{j}$. We also term $\left\{\mathcal{F}_{t}\right\}_{t\geq 0}$ the natural filtration generated by the $\mathbf{W}$.
@@
$dS_{t}^{i}=S_{t}^{i}\left(\left(r-\eta_{i}\right)dt+\sigma_{i}\Sigma_{i}d\mathbf{B}_{t}\right),$ (2.2)
where $\mathbf{B}=(\mathbf{B}_{t})_{t\in[0,T]}$ is a $d$-dimensional uncorrelated Wiener processes and $\Sigma_{i}$ is the $i$-th row of the matrix
$\Sigma=\Sigma\left(\rho_{1,2},\dots,\rho_{d-1,d}\right),$ (2.3)
which is defined as the lower Cholesky decomposition of the correlation matrix $\Gamma$, given by
@@
$\mathbf{S}_{t}=\mathbf{S}_{0}\circ\exp\left(\left(r-\boldsymbol{\eta}-\frac{\boldsymbol{\sigma}^{2}}{2}\right)t+\boldsymbol{\sigma}\circ\left(\Sigma\mathbf{B}_{t}\right)\right)$ (2.5)
with $\mathbf{v}\circ\mathbf{w}$ the Hadamard product between vectors $\mathbf{v}$ and $\mathbf{w}$, that is the element-wise product, and the exponential function in (2.5) is applied to each component of the vector $\left(r-\frac{\sigma^{2}}{2}\right)t+\sigma\circ\left(\Sigma\mathbf{B}_{t}\right)$.
Instead of assuming a single constant volatility, the UVM considers a range of possible volatilities. Specifically, the vector of volatilities $\boldsymbol{\sigma}$ is not known, but it can change over time, as a function of the time and of the underlyings, so we write $\boldsymbol{\sigma}=\boldsymbol{\sigma}\left(t,\mathbf{S}_{t}\right)=\left(\sigma_{1}\left(t,\mathbf{S}_{t}\right),\dots,\sigma_{d}\left(t,\mathbf{S}_{t}\right)\right)^{\top}$. Moreover, the explicit form of $\boldsymbol{\sigma}\left(t,\mathbf{S}_{t}\right)$ is not specified in the model, but we only know that, for each $i=1,\dots,d$ there exist two bounds for $\sigma_{i}$, that is
@@
These bounds can be estimated based on historical data by calculating the maximum and minimum values of the corresponding statistical estimators over different time windows (in this regard, see e.g. [Ruppert and Matteson, 2011]). Such a degree of uncertainty can be extended to correlations between Wiener processes. Specifically $\rho_{i,j}=\rho_{i,j}\left(t,\mathbf{S}_{t}\right)$ may change over time based on the underlying values, but the particular law is not a model datum. Instead, we only know that, for each couple of values $i=1,\dots,d-1$ and $j=i+1,\dots,d$, there exist two bounds for $\rho_{i,j}$, that is
@@
$\boldsymbol{\theta}=\boldsymbol{\theta}\left(t,\mathbf{S}_{t}\right)=\left(\sigma_{1}\left(t,\mathbf{S}_{t}\right),\dots,\sigma_{d}\left(t,\mathbf{S}_{t}\right),\rho_{1,2}\left(t,\mathbf{S}_{t}\right),\dots,\rho_{d-1,d}\left(t,\mathbf{S}_{t}\right)\right)^{\top}$
the vector of all uncertain model parameters as a function of $t$ and $\mathbf{S}_{t}$, and $\boldsymbol{\Theta}$ the set of all admissible values $\boldsymbol{\theta}$, that is, which satisfy relations (2.6), (2.7), and (2.9).
In the UVM, the pricing of derivatives is approached as an optimization problem due to the uncertainty in the volatility and correlation of the underlying asset. This model aims to find the worst-case scenario within this range to ensure robust and conservative pricing. Specifically, assuming the perspective of the contract seller , we seek to maximize the expected payoff of the derivative over all admissible scenarios for volatilities and correlation, that is over all $\boldsymbol{\theta}$ in $\Theta$. Let $V\left(t,\mathbf{S}_{t}\right)$ represent the value of the derivative at time $t$ and asset price $\mathbf{S}_{t}$ in the UVM. The pricing problem is then given by:
$V\left(t,\mathbf{S}_{t}\right)=\sup_{\boldsymbol{\theta}\in\Theta}\mathbb{E}\left[e^{-r(T-t)}\Psi\left(\mathbf{S}_{T}\right)\mid\mathbf{S}_{t}\right],$ (2.10)
where $\Psi$ is the payoff function at maturity $T$, the dynamics of $\mathbf{S}_{t}$ is given by (2.8) for $\boldsymbol{\theta}\in\Theta$, and $\mathbb{E}$ denotes the expectation under the risk-neutral measure $\mathbb{Q}$. This optimization ensures that the derivative’s price accounts for the most adverse volatility-correlation scenario, providing a conservative estimate that is robust to volatility fluctuations. By solving this maximization problem, we obtain a price that safeguards against the uncertainty in the market, making it a reliable valuation method in uncertain volatility conditions.
In addition, valuation can also be considered from the perspective of the agent who purchased the contract. In this case, the buyer would be concerned about a potential devaluation of the contract; therefore, the optimization problem becomes a minimization problem:
$V\left(t,\mathbf{S}_{t}\right)=\inf_{\boldsymbol{\theta}\in\Theta}\mathbb{E}\left[e^{-r(T-t)}\Psi\left(\mathbf{S}_{T}\right)\mid\mathbf{S}_{t}\right].$ (2.11)
In the following, for the sake of conciseness, we restrict our analysis to problem (2.10), as problem (2.11) follows a similar reasoning.
3 The GTU method
3.1 GPR-Tree for option pricing
The proposed GTU method is based on the GPR-Tree algorithm. Here we give some details about this latter algorithm and we refer the interested reader for further details to [Goudenège et al., 2020].
The GPR-Tree algorithm exploits Gaussian Process Regression and a multidimensional binomial tree to compute the price of multidimensional American option.